### v1

@@ 3 Discrete Ambient Space
To directly construct more practical and efficient a posteriori error estimates, we make use of finite-dimensional discrete ambient spaces $V$ in this section.
@@ 3.1 Practical Estimators in H(grad)
Let $\mathcal{T}_{h/2}$ be a uniform refinement of $\mathcal{T}_{h}$ and $V=V_{h/2}$ the finite element space based on the refined mesh $\mathcal{T}_{h/2}$. For simplicial meshes, $\mathcal{T}_{h/2}$ is constructed by dividing each element in $\mathcal{T}_{h}$ into four congruent children in $\mathbb{R}^{2}$ and by dividing each element in $\mathcal{T}_{h}$ into eight children (cf. [16]) in $\mathbb{R}^{3}$.
@@
Let $\mathbb{A}_{h/2}$ be the stiffness matrix of the bilinear form $a|_{V_{h/2}}$ and $S^{a}_{h/2}:V_{h/2}^{*}\rightarrow V_{h/2}$ be the operator corresponding to the inverse of the diagonal of $\mathbb{A}_{h/2}$. The following
@@
In the above iteration, $u_{n}$ converges to $u_{h/2}$ uniformly with respect to $h$ (cf. [58]). On the matrix level, $S^{a}_{h/2}$ and $S^{m}_{h/2}$ are nothing but the Jacobi and Gauss–Seidel iterators for $\mathbb{A}_{h/2}$, respectively.
@@
(3.1a) $|u_{h/2}-u_{h}|_{H^{1}(\Omega)}$ $\eqsim\langle r_{h/2},S^{a}_{h/2}r_{h/2}\rangle^{\frac{1}{2}},$
(3.1b) $|u_{h/2}-u_{h}|_{H^{1}(\Omega)}$ $\eqsim|S^{m}_{h/2}r_{h/2}|_{H^{1}(\Omega)}.$
@@
(3.3) $\frac{1}{1+\gamma_{h/2}}|u_{h/2}-u_{h}|_{H^{1}(\Omega)}\leq|u-u_{h}|_{H^{1}(\Omega)}\leq\frac{1}{1-\gamma_{h/2}}|u_{h/2}-u_{h}|_{H^{1}(\Omega)}.$
@@
(3.4a) $|u-u_{h}|_{H^{1}(\Omega)}$ $\eqsim\langle r_{h/2},S^{a}_{h/2}r_{h/2}\rangle^{\frac{1}{2}},$
(3.4b) $|u-u_{h}|_{H^{1}(\Omega)}$ $\eqsim|S^{m}_{h/2}r_{h/2}|_{H^{1}(\Omega)}.$
Replacing $V=V_{h/2}$ with a finite element space $V=V_{h,\mathcal{P}_{2}}$ based on piecewise quadratic polynomials on $\mathcal{T}_{h}$, we obtain another smoother-type a posteriori error estimate. Let $S_{h,\mathcal{P}_{2}}^{a}:V_{h,\mathcal{P}_{2}}^{*}\rightarrow V_{h,\mathcal{P}_{2}}$ and $S^{m}_{h,\mathcal{P}_{2}}:V_{h,\mathcal{P}_{2}}^{*}\rightarrow V_{h,\mathcal{P}_{2}}$ be the operator corresponding to the inverse of the diagonal and the inverse of the upper triangular part of the $\mathcal{P}_{2}$ finite element matrix on $\mathcal{T}_{h}$, respectively. Let $u_{h,\mathcal{P}_{2}}$ be the finite element solution in $V_{h,\mathcal{P}_{2}}$ and $A_{h,\mathcal{P}_{2}}:V_{h,\mathcal{P}_{2}}\rightarrow V_{h,\mathcal{P}_{2}}^{*}$ be given by
@@
Under another saturation assumption (cf. [21])
(3.5) $\|u-u_{h,\mathcal{P}_{2}}\|_{V}\leq\gamma_{\mathcal{P}_{2}}\|u-u_{h}\|_{V},\quad\gamma_{\mathcal{P}_{2}}<1,$
@@
(3.6a) $|u-u_{h}|_{H^{1}(\Omega)}$ $\eqsim\langle r_{\mathcal{P}_{2}},S^{a}_{h,\mathcal{P}_{2}}r_{\mathcal{P}_{2}}\rangle^{\frac{1}{2}},$
(3.6b) $|u-u_{h}|_{H^{1}(\Omega)}$ $\eqsim|S^{m}_{h,\mathcal{P}_{2}}r_{\mathcal{P}_{2}}|_{H^{1}(\Omega)},$
@@
Figure 1: Assignment of $\langle r_{h/2},S^{a}_{h/2}r_{h/2}\rangle$ to each element in the refinement $\mathcal{T}_{h/2}$.
Figure 2: Convergence of error estimators based on fine-coarse (left) and $\mathcal{P}_{2}$-$\mathcal{P}_{1}$ (right) layers.
Figure 3: Convergence of $\mathcal{P}_{p}$-AFEMs driven by pointwise smoother-type error estimators with effectivity ratio estimator/error shown in the parenthesis.
Figure 4: Adaptive error estimator convergence in the $L^{2}$ norm for the Poisson equation in fine-coarse layers(left) and P2-P1 layers (right).
@@
To guide the local mesh refinement, each $a_{ii}^{-1}r_{i}^{2}$ in $\langle r_{h/2},S^{a}_{h/2}r_{h/2}\rangle$ is equally subdivided and allocated to all elements in $\mathcal{T}_{h/2}$ touching the vertex $a_{h/2,i}\in\mathcal{V}_{h/2}$, see Figure 1 for example. The error estimator in (3.6a) is localized to each triangle of $\mathcal{T}_{h}$ in a similar way. Naturally, $|S^{m}_{h/2}r_{h/2}|_{H^{1}(\Omega)}$ and $|S^{m}_{h,\mathcal{P}_{2}}r_{\mathcal{P}_{2}}|_{H^{1}(\Omega)}$ are localized based on the following splitting:
@@
On the L-shaped domain $\Omega=(-1,1)^{2}\backslash([0,1)\times[-1,0))$, we test the performance of the proposed a posteriori error estimates for the FEM (2.8b). The exact solution is $u=r^{2/3}\sin(2\theta/3)$ with singularity near the re-entrant corner of $\Omega$, where $(r,\theta)$ is the polar coordinate near the origin. All quantities are computed on adaptively generated meshes by the standard AFEM with a Dörfler marking parameter $\theta=0.5$ (cf. [28, 26, 17]), which will be used in all numerical experiments. The smoother-type a posteriori error estimates as well as the implicit estimators (2.9) and (2.10) are compared to the explicit residual error estimator
@@
$\mathcal{E}_{\rm res}=\Big(\sum_{T\in\mathcal{T}_{h}}h_{T}^{2}\|f+\Delta u_{h}\|_{L^{2}(T)}^{2}+\sum_{E\text{ is an interior edge in }\mathcal{T}_{h}}h_{E}\|\llbracket\partial_{n}u_{h}\rrbracket\|^{2}_{L^{2}(E)}\Big)^{\frac{1}{2}},$
@@
where $h_{T}={\rm diam}(T)$, $h_{E}={\rm diam}(E)$, and $\llbracket\partial_{n}u_{h}\rrbracket|_{E}$ is the jump of of the normal derivative $\partial_{n}u_{h}$ across an edge $E$.
@@
We plot error curves for various AFEMs driven by different error estimators to assess convergence. It is observed from Figure 2 that the smoother-type estimators are more accurate than $\mathcal{E}_{\rm res}$. Meanwhile, they are no worse than the more expensive error estimators (2.9) and (2.10) based on local Dirichlet problems on vertex patches. In the comparison, the most efficient a posteriori estimates are based on diagonal smoothers $S_{h/2}^{a}$ and $S_{\mathcal{P}_{2}}^{a}$.
@@
Smoother-type a posteriori error estimates could be directly generalized to higher order finite elements. Let $S_{\mathcal{P}_{p}}^{a}:V_{\mathcal{P}_{p}}^{*}\rightarrow V_{\mathcal{P}_{p}}$ and $S^{m}_{\mathcal{P}_{p}}:V_{\mathcal{P}_{p}}^{*}\rightarrow V_{\mathcal{P}_{p}}$ be the operator corresponding to the inverse of the diagonal and the inverse of the upper triangular part of the $\mathcal{P}_{p}$ finite element matrix on $\mathcal{T}_{h/2}$, respectively. With $r=r_{h/2}$, the polynomial-degree-robustness ($p$-robustness) of smoother-type estimators $\langle r,S_{\mathcal{P}_{p}}^{a}r\rangle^{\frac{1}{2}}$ and $|S_{\mathcal{P}_{p}}^{a}r|_{H^{1}(\Omega)}$ has been shown in Figure 3. It is well-known that $\mathcal{E}_{\rm res}$ is not $p$-robust, i.e., $\mathcal{E}_{\rm res}/|u-u_{h}|_{H^{1}(\Omega)}$ is at least $O(p)$ grows as the finite element degree $p$ increases (cf. [46]).
@@
Compared to $\langle r_{h/2},S_{h/2}^{a}r_{h/2}\rangle^{\frac{1}{2}}$, we note that $|S_{h/2}^{a}r_{h/2}|_{H^{1}(\Omega)}$ is more flexible as it is easily adapted to other norms. For instance, the smoother-type estimator $\|S_{h/2}^{a}r_{h/2}\|_{L^{2}(\Omega)}$ is able to guide the adaptive FEM for reducing the $L^{2}$ error $\|u-u_{h}\|_{L^{2}(\Omega)}$, see Figure 4. Moreover, the a posteriori error bound $\|S_{h/2}^{a}r_{h/2}\|_{L^{2}(\Omega)}$ is sharper than the following residual error estimator

### v2

@@ 3 Discrete Ambient Space
To directly construct more practical and efficient a posteriori error estimates, we make use of finite-dimensional discrete ambient spaces $V$ arising from mesh refinement or polynomial enrichment.
@@ 3.1 Practical Estimators in H(grad)
Let $\mathcal{T}_{h/2}$ be a uniform refinement of $\mathcal{T}_{h}$ and $V=V_{h/2}$ be the finite element space based on the refined mesh $\mathcal{T}_{h/2}$. For simplicial meshes, $\mathcal{T}_{h/2}$ is constructed by dividing each element in $\mathcal{T}_{h}$ into four congruent children in $\mathbb{R}^{2}$ and by dividing each element in $\mathcal{T}_{h}$ into eight children (cf. [18]) in $\mathbb{R}^{3}$.
3.1.1 Pointwise Smoother-type Error Estimators
@@
Let $\mathbb{A}_{h/2}$ be the stiffness matrix of the bilinear form $a|_{V_{h/2}}$ and $S^{a}_{h/2}:V_{h/2}^{*}\rightarrow V_{h/2}$ be the operator corresponding to the inverse of the diagonal of $\mathbb{A}_{h/2}$. Corresponding to the stable decomposition $V_{h/2}=V_{h}+\sum_{a_{h/2,k}\in\mathring{\Omega}}{\rm span}(\phi_{h/2,k})$, the following
@@
In the above iteration, $u_{n}$ converges to $u_{h/2}$ uniformly with respect to $h$ (cf. [67]). On the matrix level, $S^{a}_{h/2}$ and $S^{m}_{h/2}$ are nothing but one step of Jacobi and Gauss–Seidel smoothing pass for $\mathbb{A}_{h/2}$, respectively. On the analytic level, $S^{a}_{h/2}$ and $S^{m}_{h/2}$ amount to solving one-dimensional local problems and are referred to as pointwise smoothers.
@@
$|u_{h/2}-u_{h}|_{H^{1}(\Omega)}$ $\eqsim\langle r_{h/2},S^{a}_{h/2}r_{h/2}\rangle^{\frac{1}{2}},$ (3.1a)
$|u_{h/2}-u_{h}|_{H^{1}(\Omega)}$ $\eqsim|S^{m}_{h/2}r_{h/2}|_{H^{1}(\Omega)}.$ (3.1b)
@@
$\frac{1}{1+\gamma_{h/2}}|u_{h/2}-u_{h}|_{H^{1}(\Omega)}\leq|u-u_{h}|_{H^{1}(\Omega)}\leq\frac{1}{1-\gamma_{h/2}}|u_{h/2}-u_{h}|_{H^{1}(\Omega)}.$ (3.3)
@@
$|u-u_{h}|_{H^{1}(\Omega)}$ $\eqsim\langle r_{h/2},S^{a}_{h/2}r_{h/2}\rangle^{\frac{1}{2}},$ (3.4a)
$|u-u_{h}|_{H^{1}(\Omega)}$ $\eqsim|S^{m}_{h/2}r_{h/2}|_{H^{1}(\Omega)}.$ (3.4b)
Alternatively, we can construct smoothers based on polynomial enrichment. Replacing $V=V_{h/2}$ with the finite element space $V=V_{h,\mathcal{P}_{2}}$ based on piecewise quadratic polynomials on $\mathcal{T}_{h}$, we obtain another smoother-type a posteriori error estimate. Let $S_{h,\mathcal{P}_{2}}^{a}:V_{h,\mathcal{P}_{2}}^{*}\rightarrow V_{h,\mathcal{P}_{2}}$ and $S^{m}_{h,\mathcal{P}_{2}}:V_{h,\mathcal{P}_{2}}^{*}\rightarrow V_{h,\mathcal{P}_{2}}$ be the operator corresponding to the inverse of the diagonal and the inverse of the upper triangular part of the $\mathcal{P}_{2}$ finite element matrix on $\mathcal{T}_{h}$, respectively. Let $u_{h,\mathcal{P}_{2}}$ be the finite element solution in $V_{h,\mathcal{P}_{2}}$ and $A_{h,\mathcal{P}_{2}}:V_{h,\mathcal{P}_{2}}\rightarrow V_{h,\mathcal{P}_{2}}^{*}$ be given by
@@
Under the following $p$-version saturation assumption (cf. [23])
$\|u-u_{h,\mathcal{P}_{2}}\|_{V}\leq\gamma_{\mathcal{P}_{2}}\|u-u_{h}\|_{V},\quad\gamma_{\mathcal{P}_{2}}<1,$ (3.5)
@@
$|u-u_{h}|_{H^{1}(\Omega)}$ $\eqsim\langle r_{\mathcal{P}_{2}},S^{a}_{h,\mathcal{P}_{2}}r_{\mathcal{P}_{2}}\rangle^{\frac{1}{2}},$ (3.6a)
$|u-u_{h}|_{H^{1}(\Omega)}$ $\eqsim|S^{m}_{h,\mathcal{P}_{2}}r_{\mathcal{P}_{2}}|_{H^{1}(\Omega)},$ (3.6b)
@@
Remark 3.1.
The implicit extrapolation technique [45] provides an efficient way to assemble a $\mathcal{P}_{2}$ smoother matrix $\mathbb{S}_{h,\mathcal{P}_{2}}$ via $\mathbb{S}_{h,\mathcal{P}_{2}}=(4/3)\mathbb{S}_{h/2}-(1/3)\mathbb{S}_{h}$, where the matrices $\mathbb{S}_{h/2}$ and $\mathbb{S}_{h}$ represent the $\mathcal{P}_{1}$ smoothers $S_{h/2}$ and $S_{h}$, respectively.
Figure 3.1: Assignment of $\langle r_{h/2},S^{a}_{h/2}r_{h/2}\rangle$ to each element in the refinement $\mathcal{T}_{h/2}$.
3.1.2 Localized Error Estimator and Adaptive Algorithm
@@
To guide the local mesh refinement, each $a_{ii}^{-1}r_{i}^{2}$ in $\langle r_{h/2},S^{a}_{h/2}r_{h/2}\rangle$ is equally subdivided (according to the degree ${\rm deg}(a_{h/2,i})$ of the vertex $a_{h/2,i}$ in $\mathcal{T}_{h/2}$) and allocated to all elements in $\mathcal{T}_{h/2}$ touching the vertex $a_{h/2,i}\in\mathcal{V}_{h/2}$, see Figure 3.1 for example. The local estimator for each $T\in\mathcal{T}_{h}$ reads
$\mathcal{E}_{T}=\left(\sum_{t\subset T,t\in\mathcal{T}_{h/2}}\sum_{a_{h/2,i}\in\bar{t}}\frac{1}{{\rm deg}(a_{h/2,i})}a_{ii}^{-1}r_{ii}^{2}\right)^{\frac{1}{2}}.$}
The error estimator in (3.6a) is localized to each triangle of $\mathcal{T}_{h}$ in a similar way. Naturally, $|S^{m}_{h/2}r_{h/2}|_{H^{1}(\Omega)}$ and $|S^{m}_{h,\mathcal{P}_{2}}r_{\mathcal{P}_{2}}|_{H^{1}(\Omega)}$ are localized based on the following splitting:
@@
Algorithm 3.1 AFEM driven by Smoother-type Error Estimator
1:Input: current mesh $\mathcal{T}_{h}$; // 1. SOLVE
2: Solve the finite element system on $\mathcal{T}_{h}$ to obtain $u_{h}$; // 2. ESTIMATE
3: Construct auxiliary fine mesh $\mathcal{T}_{h/2}$;
4: Compute fine-grid problem-specific residual $r_{h/2}\in V_{h/2}^{*}$: $\langle r_{h/2},v\rangle=(f,v)-a(u_{h},v),\quad\forall v\in V_{h/2}$;
5: Assemble the smoother $S_{h/2}$ and compute $\langle r_{h/2},S_{h/2}r_{h/2}\rangle$ or $\|S_{h/2}r_{h/2}\|$;// 3. ADAPT
6: Localize $\langle r_{h/2},S_{h/2}r_{h/2}\rangle$ or $\|S_{h/2}r_{h/2}\|$ to obtain $\{\mathcal{E}_{T}\}_{T\in\mathcal{T}_{h}}$;
7: Mark a subset $\mathcal{M}_{h}\subset\mathcal{T}_{h}$ based on Dörfler’s criterion with $\theta\in(0,1)$: $\sum_{T\in\mathcal{M}}\mathcal{E}_{T}^{2}\geq\theta^{2}\sum_{T\in\mathcal{T}_{h}}\mathcal{E}_{T}^{2}$;
8: Update $\mathcal{T}_{h}$ by refining $\mathcal{M}_{h}$ and neighboring elements based on newest vertex bisection, go to Step 1;
Algorithm 3.1 outlines an adaptive mesh refinement procedure driven by smoother-type a posteriori error estimates. It is characterized by several inexpensive operations on an auxiliary fine mesh, which avoid solving any global system. In particular, the residual evaluation, construction of smoothers, and smoothing of the residual all exhibit linear complexity. The adaptive feedback loop follows the classical AFEM framework (cf. [33, 28, 19]) based on Dörfler’s marking and newest vertex bisection rule.
In all numerical experiments, the marking parameter $\theta=0.5$ in Algorithm 3.1 will be used. Although the PDE, the domain, the finite element space $V_{h}$, the residual $r_{h}\in V_{h}^{*}$, and the estimator $\mathcal{E}_{T}$ may vary from case to case, we adhere to the general adaptive workflow outlined in Algorithm 3.1. The quantity shown in parentheses in each figure legend is the effectivity ratio, i.e., the value of estimator/error.
3.1.3 Numerical Experiments in H(grad)
@@
On the L-shaped domain $\Omega=(-1,1)^{2}\backslash([0,1)\times[-1,0))$, we test the performance of the proposed a posteriori error estimates for the FEM (2.9b). The exact solution is $u=r^{2/3}\sin(2\theta/3)$ with singularity near the re-entrant corner of $\Omega$, where $(r,\theta)$ is the polar coordinate near the origin. All quantities are computed on adaptively generated meshes by Algorithm 3.1. The smoother-type a posteriori error estimates as well as computable discrete approximations of implicit estimators (2.10) and (2.11) are compared to the explicit residual error estimator
@@
$\mathcal{E}_{\rm res}=\Big(\sum_{T\in\mathcal{T}_{h}}h_{T}^{2}\|f+\Delta u_{h}\|_{L^{2}(T)}^{2}+\sum_{E\text{ is an interior edge in }\mathcal{T}_{h}}h_{E}\|\llbracket\partial_{n}u_{h}\rrbracket\|^{2}_{L^{2}(E)}\Big)^{\frac{1}{2}},$ (3.7)
@@
where $h_{T}={\rm diam}(T)$, $h_{E}={\rm diam}(E)$, and $\llbracket\partial_{n}u_{h}\rrbracket|_{E}$ is the jump of the normal derivative $\partial_{n}u_{h}$ across an edge $E$.
Figure 3.2: Convergence of error estimators based on fine-coarse (left) and $\mathcal{P}_{2}$-$\mathcal{P}_{1}$ (right) layers.
Figure 3.3: Convergence of $\mathcal{P}_{p}$-AFEMs driven by pointwise smoother-type error estimators.
Remark 3.2 (Inhomogeneous Dirichlet boundary condition).
For inhomogeneous and piecewise polynomial boundary data $u|_{\partial\Omega}=g_{h}\in V_{h}|_{\partial\Omega}$, the finite element solution $u_{h}\in V_{h}+g_{h}$ exactly matches $u$ and $u_{h/2}$ on the boundary. As a result, $u-u_{h}\in H_{0}^{1}(\Omega)={\rm Dom}(-\Delta)$ and $u_{h/2}-u_{h}\in V_{h/2}={\rm Dom}(A_{h/2})$ are contained in the domain of $\Delta:H_{0}^{1}(\Omega)\rightarrow H^{-1}(\Omega)$ and $A_{h/2}$, respectively. In this case, the same smoothers for $\Delta$ and $A_{h/2}$ in Sections 2.2 and 3.1 are applied to the residual to obtain smoother-type a posteriori error estimates. For general boundary data $u|_{\partial\Omega}=g$, the data oscillation $\big(\sum_{{\rm edge}~E\subset\partial\Omega}h_{E}\|g-g_{h}\|_{L^{2}(E)}^{2}\big)^{\frac{1}{2}}$ is added to the error estimator, where $g_{h}\in V_{h}|_{\partial\Omega}$ is an approximation of $g$.
@@
We plot error curves for various AFEMs driven by different error estimators to assess convergence. It is observed from Figure 3.2 that the smoother-type estimators are more accurate than $\mathcal{E}_{\rm res}$. Meanwhile, they are no worse than the more expensive error estimators (2.10) and (2.11) based on local Dirichlet problems on vertex patches. In the comparison, the most efficient a posteriori estimates are based on diagonal smoothers $S_{h/2}^{a}$ and $S_{h,\mathcal{P}_{2}}^{a}$.
@@
Smoother-type a posteriori error estimates could be directly generalized to higher order finite elements. Let $S_{\mathcal{P}_{p}}^{a}:V_{\mathcal{P}_{p}}^{*}\rightarrow V_{\mathcal{P}_{p}}$ and $S^{m}_{\mathcal{P}_{p}}:V_{\mathcal{P}_{p}}^{*}\rightarrow V_{\mathcal{P}_{p}}$ be the operator corresponding to the inverse of the diagonal and the inverse of the upper triangular part of the $\mathcal{P}_{p}$ finite element matrix on $\mathcal{T}_{h/2}$, respectively. With $r=r_{h/2}$, the polynomial-degree-robustness ($p$-robustness) of smoother-type estimators $\langle r,S_{\mathcal{P}_{p}}^{a}r\rangle^{\frac{1}{2}}$ and $|S_{\mathcal{P}_{p}}^{a}r|_{H^{1}(\Omega)}$ is numerically demonstrated in Figure 3.3. In addition, Table 3.1 reports effectivity ratios of the residual estimator $\mathcal{E}_{\rm res}$ and the equilibrated estimator $\mathcal{E}_{\rm eq}$ in [36] for controlling the $\mathcal{P}_{p}$ finite element errors with $1\leq p\leq 7$. The numerical results confirm that $\mathcal{E}_{\rm res}$ is not $p$-robust: $\mathcal{E}_{\rm res}/|u-u_{h}|_{H^{1}(\Omega)}$ grows at least linearly in $p$ increases. In contrast, $\mathcal{E}_{\mathrm{eq}}$ provides sharp effectivity ratios close to 1, albeit at the cost of a more involved, implicit patchwise flux reconstruction.
Inspired by multigrid methods that perform multiple smoothing on each level, we consider the smoother-type error estimator $|S^{m,\text{iter}}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$, which corresponds to applying iter steps of Gauss–Seidel smoothing to the residual. Table 3.2 shows that raising iter from 1 to 4 yields only a modest gain in the effectivity ratio. Therefore, we adopt a single smoothing pass in all subsequent experiments.
@@
Compared to $\langle r_{h/2},S_{h/2}^{a}r_{h/2}\rangle^{\frac{1}{2}}$, we note that $|S_{h/2}^{a}r_{h/2}|_{H^{1}(\Omega)}$ is more flexible as it is easily adapted to other norms. For instance, the smoother-type estimator $\|S_{h/2}^{a}r_{h/2}\|_{L^{2}(\Omega)}$ is able to guide the adaptive FEM for reducing the $L^{2}$ error $\|u-u_{h}\|_{L^{2}(\Omega)}$, see Figure 3.4. Moreover, the a posteriori error bound $\|S_{h/2}^{a}r_{h/2}\|_{L^{2}(\Omega)}$ is sharper than the following residual error estimator
@@
Figure 3.4: Adaptive error estimator convergence in the $L^{2}$ norm for the Poisson equation in fine-coarse layers (left) and $\mathcal{P}_{2}$-$\mathcal{P}_{1}$ layers (right).
Table 3.1: Effectivity ratios of smoother-type $\langle r,S^{a}_{\mathcal{P}_{p}}r\rangle^{1/2}$ and $|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$, residual $\mathcal{E}_{\rm{res}}$, and equilibrated $\mathcal{E}_{\mathrm{eq}}$ error estimators.
Effectivity ratio Polynomial Order
$\mathcal{P}_{1}$ $\mathcal{P}_{2}$ $\mathcal{P}_{3}$ $\mathcal{P}_{4}$ $\mathcal{P}_{5}$ $\mathcal{P}_{6}$ $\mathcal{P}_{7}$
$|S^{m}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{H^{1}(\Omega)}$ 0.8632 0.9067 0.8279 0.7641 0.7183 0.7391 0.7988
$\langle r,S^{a}_{\mathcal{P}_{p}}r\rangle^{1/2}/|u-u_{h}|_{H^{1}(\Omega)}$ 0.8899 0.9184 0.8953 0.8464 0.5690 0.6166 0.7255
$|S^{a}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}/|u-u_{h}|_{H^{1}(\Omega)}$ 0.8705 0.9754 0.9665 0.9069 0.8610 0.8481 1.2495
$\mathcal{E}_{\rm{res}}/|u-u_{h}|_{H^{1}(\Omega)}$ 4.9003 7.7551 12.8526 19.5604 26.8670 34.6710 43.0182
$\mathcal{E}_{\mathrm{eq}}/|u-u_{h}|_{H^{1}(\Omega)}$ 1.0608 1.0230 1.0250 1.0365 1.0450 1.0520 1.0507
Table 3.2: Effectivity ratios of the smoother-type error estimator $|S^{m,\text{iter}}_{\mathcal{P}_{p}}r|_{H^{1}(\Omega)}$ based on $1\leq\text{iter}\leq 4$ steps of Gauss–Seidel smoothing operations.
Iter. Polynomial Order
$\mathcal{P}_{1}$ $\mathcal{P}_{2}$ $\mathcal{P}_{3}$ $\mathcal{P}_{4}$ $\mathcal{P}_{5}$ $\mathcal{P}_{6}$ $\mathcal{P}_{7}$
1 0.8632 0.9067 0.8279 0.7641 0.7183 0.7391 0.7988
2 0.8641 0.9423 0.9143 0.8688 0.8304 0.7961 0.7378
3 0.8643 0.9405 0.9574 0.9185 0.8893 0.8415 0.7833
4 0.8643 0.9480 0.9759 0.9421 0.9130 0.8627 0.8038
Remark 3.3.
The theoretical $p$-robustness of smoother-type a posteriori error estimates introduced in this work remains an open question because the stable decomposition and contraction in the analysis are not proved to be $p$-robust. Currently, only equilibrated error estimators have been proven to be $p$-robust for specific problems (cf. [21, 37, 29, 50]), owing to their carefully designed patchwise local problems and technical polynomial lifting tools.